What is the frequency of an acrobat swinging on a trapeze if the cables are approximately 2.5 m long?


1
Expert's answer
2021-12-09T09:14:39-0500

Let's first find the period of an acrobat swing:


"T=2\pi\sqrt{\dfrac{l}{g}}=2\pi\sqrt{\dfrac{2.5\ m}{9.8\ \dfrac{m}{s^2}}}=3.17\ s."

Finally, we can find the frequency of an acrobat swing as folllows:


"f=\dfrac{1}{T}=\dfrac{1}{3.17\ s}=0.32\ Hz."